# MedNeXt: Transformer-driven Scaling of ConvNets for Medical Image Segmentation

### v1

@@ Abstract
There has been exploding interest in embracing Transformer-based architectures for medical image segmentation. However, the lack of large-scale annotated medical datasets make achieving performances equivalent to those in natural images challenging. Convolutional networks, in contrast, have higher inductive biases and consequently, are easily trainable to high performance. Recently, the ConvNeXt architecture attempted to modernize the standard ConvNet by mirroring Transformer blocks. In this work, we improve upon this to design a modernized and scalable convolutional architecture customized to challenges of data-scarce medical settings. We introduce MedNeXt, a Transformer-inspired large kernel segmentation network which introduces – 1) A fully ConvNeXt 3D Encoder-Decoder Network for medical image segmentation, 2) Residual ConvNeXt up and downsampling blocks to preserve semantic richness across scales, 3) A novel technique to iteratively increase kernel sizes by upsampling small kernel networks, to prevent performance saturation on limited medical data, 4) Compound scaling at multiple levels (depth, width, kernel size) of MedNeXt. This leads to state-of-the-art performance on 4 tasks on CT and MRI modalities and varying dataset sizes, representing a modernized deep architecture for medical image segmentation. Our code is available here: [to be added soon]
@@ 1 Introduction
Transformers [30, 7, 21] have seen wide-scale adoption in medical image segmentation as either components of hybrid architectures [3, 9, 33, 2, 8, 31] or standalone techniques [34, 25, 15] for state-of-the-art performance. The ability to learn long-range spatial dependencies is one of the major advantages of the Transformer architecture in visual tasks. However, Transformers are plagued by the necessity of large annotated datasets to maximize performance benefits owing to their limited inductive bias. While such datasets are common to natural images (ImageNet-1k [6], ImageNet-21k [26]), medical image datasets usually suffer from the lack of abundant high quality annotations [19]. To retain the inherent inductive bias of convolutions while taking advantage of architectural improvements of Transformers, the ConvNeXt [22] was recently introduced to re-establish the competitive performance of convolutional networks for natural images. The ConvNeXt architecture uses an inverted bottleneck mirroring that of Transformers, composed of a depthwise layer, an expansion layer and a contraction layer (Sec. 2.1), in addition to large depthwise kernels to replicate long-range representation learning. The authors paired large kernel ConvNeXt networks with enormous datasets to outperform erstwhile state-of-the-art Transformer-based networks. In contrast, the VGGNet [28] approach of stacking small kernels continues to be the predominant technique for designing ConvNets in medical image segmentation. Out-of-the-box data-efficient solutions such as nnUNet [13], using variants of a standard UNet [5], have still remained effective across a wide range of tasks.
@@
The ConvNeXt architecture marries the long-range spatial representation learning capabilities of Vision [7] and Swin Transformers [21] with the inherent inductive bias of ConvNets. Additionally, the inverted bottleneck design allows us to scale width (increase channels) while not being affected by kernel sizes. Effective usage in medical image segmentation would allow benefits from – 1) learning long-range spatial dependencies via large kernels, 2) less intuitively, simultaneously scaling multiple network levels. To achieve this would require techniques to combat the tendency of large networks to overfit on limited training data. Despite this, there have been recent attempts to introduce large kernel techniques to the medical vision domain. In [18], a large kernel 3D-UNet [5] was used by decomposing the kernel into depthwise and depthwise dilated kernels for improved performance in organ and brain tumor segmentation – exploring kernel scaling, while using constant number of layers and channels. The ConvNeXt architecture itself was utilized in 3D-UX-Net [17], where the Transformer of SwinUNETR [8] was replaced with ConvNeXt blocks for high performance on multiple segmentation tasks. However, 3D-UX-Net only uses these blocks partially in a standard convolutional encoder, limiting their possible benefits.
@@ 2.1 Fully ConvNeXt 3D Segmentation Architecture
In prior work, ConvNeXt [22] distilled architectural insights from Vision Transformers [7] and Swin Transformers [21] into a convolutional architecture. The ConvNeXt block inherited a number of significant design choices from Transformers, designed to limit computation costs while increasing receptive field width to learn global features, which demonstrated performance improvements over standard ResNets [10]. In this work, we leverage these strengths by adopting the general design of ConvNeXt as the building block in a 3D-UNet-like [5] macro architecture to obtain the MedNeXt. We extend these blocks to up and downsampling layers as well (Sec. 2.2), resulting in the first fully ConvNeXt architecture for medical image segmentation. The macro architecture is illustrated in Figure 1(a). MedNeXt blocks (similar to ConvNeXt blocks) have 3-layers mirroring a Transformer block and are described for a $C$-channel input as follows:
@@ 2.3 UpKern: Large Kernel Convolutions without Saturation
Large convolution kernels approximate the large attention windows in Transformers, but remain prone to performance saturation. ConvNeXt architectures in classification of natural images, despite the benefit of large datasets such as ImageNet-1k and ImageNet-21k, are seen to saturate at kernels of size $7\times 7$ [22]. Medical image segmentation tasks have significantly less data and performance saturation can be a problem in large kernel networks. To propose a solution, we borrow inspiration from Swin Transformer V2 [20] where a large-attention-window network is initialized with another network trained with a smaller attention window. Specifically, Swin Transformers use a bias matrix $\hat{B}\in\mathbb{R}^{(2M-1)\times(2M-1)}$ to store learnt relative positional embeddings, where $M$ is the number of patches in an attention window. On increasing the window size, $M$ increases and necessitates a larger $\hat{B}$. The authors proposed spatially interpolating an existing bias matrix to the larger size as a pretraining step, instead of training from scratch, which demonstrated improved performance. We propose a similar approach but customized to convolutions kernels, as seen in Figure 1(b), to overcome performance saturation. UpKern allows us to iteratively increase kernel size by initializing a large kernel network with a compatible pretrained small kernel network by trilinearly upsampling convolutional kernels (represented as tensors) of incompatible size. All other layers with identical tensor sizes (including normalization layers) are initialized by copying the unchanged pretrained weights. This leads to a simple but effective initialization technique for MedNeXt which helps large kernel networks overcome performance saturation in the comparatively limited data scenarios common to medical image segmentation.
@@ 2.4 Compound Scaling of Depth, Width and Receptive Field
Compound scaling [29] is the idea that simultaneous scaling on multiple levels (depth, width, receptive field, resolution etc) offers benefits beyond that of scaling at one single level. The computational requirements of indefinitely scaling kernel sizes in 3D networks quickly becomes prohibitive and leads us to investigate simultaneous scaling at different levels. Keeping with Figure 1(a), our scaling is tested for block count ($B$), expansion ratio ($R$) and kernel size ($k$) – corresponding to depth, width and receptive field size. We use 4 model configurations of the MedNeXt to do so, as detailed in Table 1 (Left). The basic functional design (MedNeXt-S) uses number of channels ($C$) as 32, $R=2$ and $B=2$. Further variants increase on just $R$ (MedNeXt-B) or both $R$ and $B$ (MedNeXt-M). The largest 70-MedNext-block architecture uses high values of both $R$ and $B$ (MedNeXt-L) and is used to demonstrate the ability of MedNeXt to be significantly scaled depthwise (even at standard kernel sizes). We further explore large kernel sizes and experiment with $k=\{3,5\}$ for each configuration, to maximize performance via compound scaling of the MedNeXt architecture.
@@ 4 Results and Discussion
$\blacksquare$: Better than (or equal to) top baseline    $\blacksquare$: Better than kernel:3 counterpart

### v5

@@ Abstract
There has been exploding interest in embracing Transformer-based architectures for medical image segmentation. However, the lack of large-scale annotated medical datasets make achieving performances equivalent to those in natural images challenging. Convolutional networks, in contrast, have higher inductive biases and consequently, are easily trainable to high performance. Recently, the ConvNeXt architecture attempted to modernize the standard ConvNet by mirroring Transformer blocks. In this work, we improve upon this to design a modernized and scalable convolutional architecture customized to challenges of data-scarce medical settings. We introduce MedNeXt, a Transformer-inspired large kernel segmentation network which introduces – 1) A fully ConvNeXt 3D Encoder-Decoder Network for medical image segmentation, 2) Residual ConvNeXt up and downsampling blocks to preserve semantic richness across scales, 3) A novel technique to iteratively increase kernel sizes by upsampling small kernel networks, to prevent performance saturation on limited medical data, 4) Compound scaling at multiple levels (depth, width, kernel size) of MedNeXt. This leads to state-of-the-art performance on 4 tasks on CT and MRI modalities and varying dataset sizes, representing a modernized deep architecture for medical image segmentation. Our code is made publicly available at: https://github.com/MIC-DKFZ/MedNeXt.
@@ 1 Introduction
Transformers [30, 7, 21] have seen wide-scale adoption in medical image segmentation as either components of hybrid architectures [3, 9, 33, 2, 8, 31] or standalone techniques [34, 25, 15] for state-of-the-art performance. The ability to learn long-range spatial dependencies is one of the major advantages of the Transformer architecture in visual tasks. However, Transformers are plagued by the necessity of large annotated datasets to maximize performance benefits owing to their limited inductive bias. While such datasets are common to natural images (ImageNet-1k [6], ImageNet-21k [26]), medical image datasets usually suffer from the lack of abundant high quality annotations [19]. To retain the inherent inductive bias of convolutions while taking advantage of architectural improvements of Transformers, the ConvNeXt [22] was recently introduced to re-establish the competitive performance of convolutional networks for natural images. The ConvNeXt architecture uses an inverted bottleneck mirroring that of Transformers, composed of a depthwise layer, an expansion layer and a contraction layer (Sec. 2.1), in addition to large depthwise kernels to replicate their scalability and long-range representation learning. The authors paired large kernel ConvNeXt networks with enormous datasets to outperform erstwhile state-of-the-art Transformer-based networks. In contrast, the VGGNet [28] approach of stacking small kernels continues to be the predominant technique for designing ConvNets in medical image segmentation. Out-of-the-box data-efficient solutions such as nnUNet [13], using variants of a standard UNet [5], have still remained effective across a wide range of tasks.
@@
The ConvNeXt architecture marries the scalability and long-range spatial representation learning capabilities of Vision [7] and Swin Transformers [21] with the inherent inductive bias of ConvNets. Additionally, the inverted bottleneck design allows us to scale width (increase channels) while not being affected by kernel sizes. Effective usage in medical image segmentation would allow benefits from – 1) learning long-range spatial dependencies via large kernels, 2) less intuitively, simultaneously scaling multiple network levels. To achieve this would require techniques to combat the tendency of large networks to overfit on limited training data. Despite this, there have been recent attempts to introduce large kernel techniques to the medical vision domain. In [18], a large kernel 3D-UNet [5] was used by decomposing the kernel into depthwise and depthwise dilated kernels for improved performance in organ and brain tumor segmentation – exploring kernel scaling, while using constant number of layers and channels. The ConvNeXt architecture itself was utilized in 3D-UX-Net [17], where the Transformer of SwinUNETR [8] was replaced with ConvNeXt blocks for high performance on multiple segmentation tasks. However, 3D-UX-Net only uses these blocks partially in a standard convolutional encoder, limiting their possible benefits.
@@ 2.1 Fully ConvNeXt 3D Segmentation Architecture
In prior work, ConvNeXt [22] distilled architectural insights from Vision Transformers [7] and Swin Transformers [21] into a convolutional architecture. The ConvNeXt block inherited a number of significant design choices from Transformers, designed to limit computation costs while scaling the network, which demonstrated performance improvements over standard ResNets [10]. In this work, we leverage these strengths by adopting the general design of ConvNeXt as the building block in a 3D-UNet-like [5] macro architecture to obtain the MedNeXt. We extend these blocks to up and downsampling layers as well (Sec. 2.2), resulting in the first fully ConvNeXt architecture for medical image segmentation. The macro architecture is illustrated in Figure 1(a). MedNeXt blocks (similar to ConvNeXt blocks) have 3-layers mirroring a Transformer block and are described for a $C$-channel input as follows:
@@ 2.3 UpKern: Large Kernel Convolutions without Saturation
Large convolution kernels approximate the large attention windows in Transformers, but remain prone to performance saturation. ConvNeXt architectures in classification of natural images, despite the benefit of large datasets such as ImageNet-1k and ImageNet-21k, are seen to saturate at kernels of size $7\times 7\times 7$ [22]. Medical image segmentation tasks have significantly less data and performance saturation can be a problem in large kernel networks. To propose a solution, we borrow inspiration from Swin Transformer V2 [20] where a large-attention-window network is initialized with another network trained with a smaller attention window. Specifically, Swin Transformers use a bias matrix $\hat{B}\in\mathbb{R}^{(2M-1)\times(2M-1)}$ to store learnt relative positional embeddings, where $M$ is the number of patches in an attention window. On increasing the window size, $M$ increases and necessitates a larger $\hat{B}$. The authors proposed spatially interpolating an existing bias matrix to the larger size as a pretraining step, instead of training from scratch, which demonstrated improved performance. We propose a similar approach but customized to convolutions kernels, as seen in Figure 1(b), to overcome performance saturation. UpKern allows us to iteratively increase kernel size by initializing a large kernel network with a compatible pretrained small kernel network by trilinearly upsampling convolutional kernels (represented as tensors) of incompatible size. All other layers with identical tensor sizes (including normalization layers) are initialized by copying the unchanged pretrained weights. This leads to a simple but effective initialization technique for MedNeXt which helps large kernel networks overcome performance saturation in the comparatively limited data scenarios common to medical image segmentation.
@@ 2.4 Compound Scaling of Depth, Width and Receptive Field
Compound scaling [29] is the idea that simultaneous scaling on multiple levels (depth, width, receptive field, resolution etc) offers benefits beyond that of scaling at one single level. The computational requirements of indefinitely scaling kernel sizes in 3D networks quickly becomes prohibitive and leads us to investigate simultaneous scaling at different levels. Keeping with Figure 1(a), our scaling is tested for block count ($B$), expansion ratio ($R$) and kernel size ($k$) – corresponding to depth, width and receptive field size. We use 4 model configurations of the MedNeXt to do so, as detailed in Table 1 (Left). The basic functional design (MedNeXt-S) uses number of channels ($C$) as 32, $R=2$ and $B=2$. Further variants increase on just $R$ (MedNeXt-B) or both $R$ and $B$ (MedNeXt-M). The largest 62-MedNext-block architecture uses high values of both $R$ and $B$ (MedNeXt-L) and is used to demonstrate the ability of MedNeXt to be significantly scaled depthwise (even at standard kernel sizes). We further explore large kernel sizes and experiment with $k=\{3,5\}$ for each configuration, to maximize performance via compound scaling of the MedNeXt architecture.
@@ 4 Results and Discussion
val (bold): better than or equal to ($\geq$) top baseline
val (underline): better than (>>) kernel: $3$ of same configuration